OpenStudy (anonymous):

Find all exact values of B when 0360 . 8Sin (3B)-4=2 please help

OpenStudy (anonymous):

I'm going to write 0<B<360 as 0<B<2pi instead because that is easier for me. 8Sin[3B]-4=2 8Sin[3B]=6 Sin[3B]=3/4 3B=ArcSin[3/4] B=ArcSin[3/4]/3

OpenStudy (anonymous):

I would get 16.2 as a decimal answer. But why does it ask to get all values from zero to 360? how do you find that?

OpenStudy (anonymous):

0 to 360 is a full circle. Because the answer is positive, the degrees are in both the first and 2nd quadrants. ASTC: We are A and S. Therefore, the answers are 16.2 and 180-16.2

OpenStudy (anonymous):

So when it is asking find all values, it is asking for the reference angles, correct?